OpenStudy (madeline4giles852):

The general form of the equation of a circle is x2 + y2 + 8x + 22y + 37 = 0. The equation of this circle in standard form is HELP WILL MEDAL

OpenStudy (thereisnousername):

The standard form equation of a circle is a way to express the definition of a circle on the coordinate plane.h and k are the x and y coordinates of the center of the circle so the standard form formula is: (h+k)^2+(y-k)^2=r^2 so your answer is: (x+y)^2+(8x+22y)+37 = 0

OpenStudy (madeline4giles852):

thanks :)))

OpenStudy (mathstudent55):

The standard form of the equation of a circle is \((x - h)^2 + (y - k)^2 = r^2\) where point \((h, k)\) is the center of the circle, and \(r\) is the radius. To get your equation in that form, you need to complete the square for both x and y: \(x^2 + y^2 + 8x + 22y + 37 = 0\) \(x^2 + 8x + y^2 + 22y = - 37\) \(x^2 + 8x + 16 + y^2 + 22y + 121 = - 37 + 16 + 121\) \((x + 4)^2 + (y + 11)^2 = 100\) \((x + 4)^2 + (y + 11)^2 = 10^2\)